One mole of an ideal monatomic gas is initially at a temperature of 274 K. Find the final temperature of the gas if 3260 J of heat are added to it and it does 712 J of work?

solve step by step

answer should be T = 478.4 K
Answer #1

From the first law of thermodynamics

delQ= U+ W

del Q = n CV ( Tf- Ti) + W

3260 J =n * 3/2 *R (Tf- 274) + 712 J

3260 J = 1 (1.5) (8.314) ( Tf - 274) +712 J

3260 J = 12.471 Tf - 3417.054 +712

Tf = 478.4 K